OpenStudy (anonymous):

The equal sides of an isosceles triange are 20 in long. Determine the length of the base if the area is to be maximum. This is a min-max problem use optimization and please show all working...thank you :)

OpenStudy (anonymous):

So i knwo area =1/2 b*h so we can say A=1/2x*y-maximum

OpenStudy (anonymous):

there is another way: Area of triangle is also equal to = 0.5 * the length of 2 sides * sin (angle formed by 2 sides). So Max(A) = 0.5 * 20 * 20 * sin (angle). To maximize the area sin(angle) = 1. Then you know you have a right isosceles triangle. That should give you the clue to find the base

OpenStudy (anonymous):

huh? :S

OpenStudy (anonymous):

Are you familiar with the triangle formula A=0.5 * the length of 2 sides * sin (angle formed by 2 sides)?

OpenStudy (anonymous):

no